You may want to reference (Pages 281 - 283) Section 8.2 while completing this problem. A sample of methane (CH4) has a volume of 28 mL at a pressure of 0.86 atm. What is the final volume, in milliliters, of the gas at each of the following pressures, if there is no change in temperature and amount of gas?

Part A

Calculate the final volume for the sample at 0.29 atm.

Express your answer using two significant figures.

V

V

=

  

  mL  

SubmitRequest Answer

Part B

Calculate the final volume for the sample at 2.20 atm.

Express your answer using two significant figures.

V

V

=

  

  mL  

SubmitRequest Answer

Part C

Calculate the final volume for the sample at 2600 mmHg.

Express your answer using two significant figures.

V

V

=
  mL  

SubmitRequest Answer

Part D

Calculate the final volume for the sample at 73.0 torr.

Express your answer using two significant figures.

V

V

=
  mL  

SubmitRequest Answer
Answer #1

A)

Given:

Vi = 28 mL

Pi = 0.86 atm

Pf = 0.29 atm

use:

Pi*Vi = Pf*Vf

0.86 atm * 28 mL = 0.29 atm * Vf

Vf = 83 mL

Answer: 83 mL

B)

Given:

Vi = 28 mL

Pi = 0.86 atm

Pf = 2.20 atm

use:

Pi*Vi = Pf*Vf

0.86 atm * 28 mL = 2.2 atm * Vf

Vf = 11 mL

Answer: 11 mL

C)

Given:

Vi = 28 mL

Pi = 0.86 atm

Pf = 2600.0 mm Hg

= (2600.0/760) atm

= 3.4211 atm

use:

Pi*Vi = Pf*Vf

0.86 atm * 28 mL = 3.4211 atm * Vf

Vf = 7.0 mL

Answer: 7.0 mL

D)

Given:

Vi = 28 mL

Pi = 0.86 atm

Pf = 73.0 torr

= (73.0/760) atm

= 0.0961 atm

use:

Pi*Vi = Pf*Vf

0.86 atm * 28 mL = 0.0961 atm * Vf

Vf = 251 mL

Answer: 2.5*10^2 mL